OpenStudy (anonymous):

A Doctor helped 320 women deliver their babies last year. Sixteen of these women had twins.What percent of the doctor's patient had twins. Its urgent thanks

OpenStudy (anonymous):

5%

OpenStudy (anonymous):

Can you please do the steps how you get 5% so as when ever I come across such questions again I will be able to do it myself

OpenStudy (anonymous):

Answer: = 1/20 = 1 ÷ 20 = 0.05 0.05 x 100 = 5%

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Answer: = 1/20 = 1 ÷ 20 = 0.05 0.05 x 100 = 5%

OpenStudy (anonymous):

What about this question: A Community College had its budget slashed by $3,101,000. This represents a cut of 7% of its total budget from last year .What was last year's budget?